OpenStudy (anonymous):

a ball is dropped from a tower. when it lands, its velocity is 30 m/s g=10 m/s^2. how long does it take for the ball to hit the ground?

OpenStudy (abhisar):

Here, Final velocity = 30 m/s In case of dropping, initial velocity = 0 m/s acceleration = 10 m/s We have to find time taken, we can use the equation v = u+at => 30=0+10\(\sf \times t\) =>t = 3 seconds

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

Yes thank you so much! I'm so thankful....

OpenStudy (abhisar):

My pleasure :)

OpenStudy (anonymous):

you divided the velocity with the g=10 m/s^2 right? to get 3 seconds?

OpenStudy (abhisar):

yes

OpenStudy (anonymous):

okay! thanks I still have more questions :) i need to post

OpenStudy (abhisar):

Ok go ahead, post it as a new one :)